# Vision foundation models: can they be applied to astrophysics data?

### v3

@@ 2.3 UMAP Illustration
(a) GMNIST, DINOv2
(b) GMNIST, MSN
(c) GMNIST, MAE
(d) GMNIST, ResNet50
(e) RGZ, DINOv2
(f) RGZ, MSN
(g) RGZ, MAE
(h) RGZ, ResNet50
Figure 2: GMNIST (upper row) and RGZ (lower row) UMAP for features extracted using DINOv2, MSN, MAE, and Resnet50.
@@
The UMAP representation of the DINO and MSN feature spaces for GMNIST is quite similar. Classes SR and U, representing smooth-and-round and unbarred spiral galaxies, are in distinct clusters, although MSN features show more overlap between the two than DINO’s. The majority of SR and U galaxy images are separated by ResNet50 as well, although small clusters of these find themselves in other areas. Galaxies that appear predominantly elliptical – smooth-and-cigar-shaped (SC) and edge-on-disk (E) – share similar latent space embeddings; clustering is most distinct for DINOv2.
Latent space grouping according to class is far less evident with the RGZ data. The best visible separation is seen in ResNet50. The ResNet model we used was trained with the goal of image classification, as opposed to DINO and MSN, which seek to reconcile local and global image characteristics, and MAE, whose objective is image reconstruction.
@@ 2.4 Downstream Task Study Methodology
For image classification, we used the models ‘out-of-the-box’ by attaching a single-layer linear classifier head to the pre-trained foundation model backbone. Only the classifier was allowed to train, so the backbone is considered frozen in that the weights and biases of the backbone stay fixed. Full fine-tuning of the system by also allowing the backbone parameters to update is known to improve performance, but corresponding compute time is over a hundred times more. As we are investigating model selection, we do not compare with full fine-tuning or hyperparameter optimization. We do compare results with fully supervised classifiers; in this case, both the backbone network used as well as the classifier head were allowed to train. Except for supervised training, all tasks were able to be performed on a NVIDIA GeForce RTX laptop GPU. Training and evaluation took between 10 seconds (10% of training data) to 45 seconds (all training data) per run. Supervised training took from 30 minutes (10% of training data) to 4 hours (all training data) on two NVIDIA TITAN Xp GPUs.
Source detection was done in a similar fashion, using Faster-RCNN [37]. The implementation of Faster-RCNN is slightly different for Vision Transformers and CNNs. For a Vision Transformer backbone, a simple feature pyramid network (FPN) based on only the output of the last large-stride feature map of the backbone is used [26]. The Faster-RCNN implementation using ResNet also uses a FPN, but one that is heirarchical, acting on feature maps from different convolutional blocks in the ResNet rather than just the last one.
After the backbone and FPN, object detection is done by applying a sliding window of the region proposal network (RPN) that predicts whether an object is present or not. These proposals are pooled (RoI pooling) and finally bounding boxes and classes are predicted. We considered the FPN, RPN and RoI layers to comprise the detection head in this situation, as these components are what is attached on top of the backbone network.
We again considered the case of a frozen backbone and trainable detection head and compare this to networks trained in a fully supervised fashion. For this source detection task, we also compared with fine-tuning the backbone, as compute time was on the same order of magnitude whether the backbone was frozen or fine-tuned (unlike for classification). Training time on two NVIDIA TITAN Xp GPUs was between approximately 6 to 10 hours depending on the architecture and whether the backbone was frozen or fine-tuned. Training from scratch took 5 to 11 hours.
@@ 3 Galaxy Morphology Classification
Galaxy morphology classification is a common task in both optical and radio astronomy, and large labeled datasets have been produced thanks to Galaxy Zoo projects [51, 49]. Previous deep learning approaches in optical classification have relied heavily on CNNs trained from scratch (e.g. [14, 30, 34, 44]). A few works have fine-tuned ImageNet-trained models, such as [18]’s DenseNet121 and [21]’s EfficientNetB5. Some works have investigated combining the advantages of ViT’s attention mechanism with CNNs, such as [50], [11] and [10]. However, application of pure ViTs has been limited and resulted in lower performance than CNNs [25], although [22] reports fine-tuning to be better than training from scratch. Radio galaxy classification work has been more limited due to lack of large datasets and a well-defined labeling scheme, and so far only employs CNNs; [4] and [32] provide comprehensive reviews of the topic. Due to the difficult nature of the data, traditional ML techniques, feature engineering, and finding augmentations specific to radio images, are also active fields of study, while some seek to enhance training datasets with synthetic images using generative deep learning.
We demonstrate the galaxy morphology classification task with both optical and radio datasets labeled courtesy of the Galaxy Zoo project. Twenty percent of RGZ images contain two labeled galaxies, and four percent contain more than two; these images were all excluded from the classification dataset, reducing the total number of samples to 4692 training images and 1189 test images. Therefore the training data available for RGZ was just over half that available for GMNIST, which was split with 80% of its 10K images in the training set. The images in both datasets were already normalized, so the only data transformation we performed was to scale the images up to the same size (224 $\times$× 224 pixels) from their native resolutions (GMNIST: 64 $\times$× 64, RGZ: 132 $\times$× 132).
Hyperparameters for the classifier training were similar for both datasets, using a learning rate of 0.0005 and a batch size of 64. Classifiers were trained for 100 epochs on GMNIST and 200 on RGZ. Class weights inversely proportional to the number of samples present in the training dataset were used in the cross-entropy loss function used to classify RGZ. To illustrate the performance differences in low- and high-label regimes, the number of samples used for training was varied at 10%, 30%, 50% and 100%, while the test set was kept the same. For comparison, we trained a ViT-Base model with patch size 16 from scratch in a fully-supervised manner, as well as ResNets with both 50 and 18 layers. The best-performing supervised model in both cases was ResNet50.
(a) GMNIST
(b) RGZ
@@
Figure 3 shows classification F1 score as a function of training label percentage. While results differ according to dataset, some trends are in common. Performance generally improved when the number of training labels increased, MAE was the worst performer overall, and the smaller ResNet18 did not classify as well as ResNet50.
@@
DINOv2 and MSN both performed better than the ResNets, although ResNet50 improved the most when trained on 100% of the data. One might explain the superior performance of DINO and MSN by the number of parameters of their ViT-Base architectures, were it not for MAE which is also ViT-Base. Unlike DINO and MSN, which use student-teacher paradigms to reconcile partially masked or locally cropped views of images, MAE’s objective is to reconstruct an entire image from a few patches. We speculate that MAE retains too much information that is irrelevant to classification; because large portions of GMNIST images consist of empty space, this might overly contribute to the learned representations.
Classification was more difficult with Radio Galaxy Zoo, with F1 scores of less than 0.6 even when using all available training labels. This could be because more of the image is dominated by noise, and less by the galaxy to be classified. The labeling schematic could be another reason for low scores, as discussed in Section 2.1. Only MSN out-performed supervised training, and only for a low percentage of training labels.
The poor performance on radio galaxy classification is not consistent with results from previous works such as [41] and [24], which have shown much higher F1 scores (up to 0.94 for DINOv2) achieved on the MiraBest dataset [35] using pre-trained foundation models. MiraBest is designed for binary classification in the Fanarhoff-Riley scheme, which could be roughly described as a single Gaussian component with one peak versus a single component with two peaks.
@@
Unlike GMNIST, RGZ images are reconstructions which contain noise with strong inter-pixel correlations on the scale of the synthesized beam. These statistical differences between ImageNet and radio images suggest that a well-performing model may require knowledge of a very specific set of features; a true case of distribution shift. While the out-of-the-box application of foundation models to GMNIST images performed well, it is not sufficient for classification on RGZ where it offers little to no advantage over supervised training.
@@ 4 Source detection
Source detection in both optical and radio images is largely performed analytically. Work using deep learning to replace the time-consuming analytic pipelines has been done by [52, 15, 2, 38], mainly through application of Faster- or Mask-RCNN. [40] provides an overview of both CNN and ViT-based methods, including YOLO. More recent results based on simulated radio data include [42] and [8], which both surpass the capabilities of analytic source detection.
Source detection was performed on radio continuum datasets RGZ, which mostly contains single galaxies labeled by morphology, and MGCLS, where a single image contains from 10-100 individual compact sources. MGCLS images may also contain larger sources which are not labeled.
@@
Table 4: Source detection results
@@
As was done for classification, images were upscaled from their native resolution. For RGZ, scaling to $\times×224$ meant that the average single object size became $\times×55$ pixels, while the 25th percentile were $\times×24$ pixels. MGCLS crops were doubled in size from $\times×256$. This was done so that the average compact source occupied $\times×16$ pixels, and the 25th percentile $\times×14$ pixels. The size of the sources relative to the transformer patch size is important; sources unable to occupy most of a patch were difficult to detect. This was proven empirically – performance on detection with MGCLS increased by more than 10% when the image size doubled.
@@
We chose to use Faster-RCNN to compare performance on object detection. Our first experiment kept the backbones frozen, allowing only the detection head layers (FPN, RPN and ROI) to train. Due to differences in architecture this resulted in a different number of trainable parameters: 21M for ViT-Det, 17.8M for Resnet50 with FPN, and 11M for Resnet18. Because of the structure of the FPN for ViT-Det, for this task the DINO backbone used was DINOv1, which had available pre-trained weights for the architecture ViT-Base with 16x16 patch size (DINOv2 uses 14x14). In the second experiment we allowed the weights and biases of the backbone networks to update as well (fine-tuning).
We kept the default training data augmentations: a randomly applied blur, contrast adjustment, and color jitter. Different learning rates, varied in order to ensure a smooth decay of the training loss, were required depending on the dataset and network. Training was for 100 epochs, with a batch size of 16. Table 4 reports the mean average precision at an intersection-over-union (IOU) threshold of 50%, known as the mAP@50.
@@
Results showed that MSN and DINO are good foundation models for this task. They outperformed training from scratch, even when the backbone was frozen, except for MSN on MGCLS. Although the ResNet models did not outperform training from scratch when used as a frozen backbone, they equaled or exceeded that performance when fine-tuned. Detection on galaxies in the RGZ dataset appears to be a task particularly suited to ResNets, while Vision Transformers do very well at detecting compact sources in MGCLS. Generally, detection was very good on MGCLS for networks larger than ResNet18. Possible reasons for the higher mAP@50 include the single category (compact source) and the tens of sources present in the images compared to the 1-6 sources in RGZ. Additionally, bounding boxes in RGZ can be quite large, including a lot of noise and making them harder to predict accurately.
@@
These results show that transfer learning, starting with the pre-trained weights of a foundation model, is a good practice, as performance from even a frozen backbone was not worse than training from scratch, and fine-tuning improved it. It is important to ensure that properties of the architecture, such as patch size in the case of ViTs, are complementary to the data being used for the downstream task.

### v4

@@ 2.3 UMAP Illustration
Figure 2: GMNIST (upper row) and RGZ (lower row) UMAP for features extracted using DINOv2, MSN, MAE, SigLIP, AM-RADIO, and ResNet-50 (from left to right).
@@
As shown in Figure 2, the UMAP representation of the DINO, MSN, SigLIP, and AM-RADIO feature spaces for GMNIST is quite similar. Classes SR and U, representing smooth-and-round and unbarred spiral galaxies, are in distinct clusters, although MSN features show more overlap between the two than the rest. The majority of SR and U galaxy images are separated by MAE and ResNet-50 as well, although small clusters of these find themselves in other areas. Galaxies that appear predominantly elliptical – smooth-and-cigar-shaped (SC) and edge-on-disk (E) – share similar latent space embeddings; clustering is most distinct for DINOv2.
Latent space grouping according to class is far less evident with the RGZ data. The best visible separation is seen in ResNet-50 and AM-RADIO. The ResNet model we used was trained with the goal of image classification, as opposed to DINO and MSN, which seek to reconcile local and global image characteristics, and MAE, whose objective is image reconstruction.
@@ 2.4 Downstream Task Study Methodology
We examine three approaches to finding appropriate models and datasets for tasks. The first is perhaps the simplest: fine-tuning the model to the given dataset. The second, which is often closely related the other two, is to choose the model based on the downstream task. The third is to adapt the data to suit the given model.
Fine-tuning a foundation model to perform a task on a particular dataset is commonly known as transfer learning. This involves unfreezing parameters of the backbone network and allowing those gradients to update. It can be done layer-by-layer or block-by-block, in order to achieve maximum performance with the minimum fine-tuned parameters, or by unfreezing the entire backbone. In this paradigm, the weights and biases of the model itself are optimized for a given dataset and downstream task. In this work, we will consider fine-tuning for both image classification and source detection.
Related to transfer learning is the idea of adapting the model to the downstream task. Both this method and full fine-tuning involve fitting a projector head to the foundation model backbone, to enable it to carry out the DST. However, one can choose to only update the parameters in the head rather than the entire head plus backbone configuration. This significantly reduces computational cost. Choosing a head architecture that is able to filter task-relevant features from the backbone’s latent space should perform well for results for minimal compute.
For this work, the projector head used for image classification was a single-layer linear classifier. Source detection was done using Faster-RCNN (Ren et al. 2017), whose projector head is implemented slightly different for Vision Transformers and CNNs. For a Vision Transformer backbone, a simple feature pyramid network (FPN) based on only the output of the last large-stride feature map of the backbone is used (Li et al. 2022b). The Faster-RCNN implementation using ResNet also uses a FPN, but one that is heirarchical, acting on feature maps from different convolutional blocks in the ResNet rather than just the last one. After the backbone and FPN, object detection is done by applying a sliding window of the region proposal network (RPN) that predicts whether an object is present or not. These proposals are pooled (RoI pooling) and finally bounding boxes and classes are predicted. We considered the FPN, RPN and RoI layers to comprise the detection head in this situation, as these components are what is attached on top of the backbone network.
The first two approaches are the most familiar to machine learning practitioners; the last will be more intuitive to astrophysicists. Adapting the data to suit the chosen model simply means finding the optimal data pre-processing method, taking into account the characteristics of both the model and the data. In this paper, we closely examine several ways in which this can affect image classification.
@@ 3 Galaxy morphology classification
Galaxy morphology classification is a common task in both optical and radio astronomy, and large labeled datasets have been produced thanks to Galaxy Zoo projects (Walmsley et al. 2022; Wong et al. 2024). Previous deep learning approaches in optical classification have relied heavily on CNNs trained from scratch (Zhu et al. 2019; Maslej-Krešňáková et al. 2021; Pandya et al. 2023; Urechiatu & Frincu 2024, e.g.). A few works have fine-tuned ImageNet-trained models, such as Hui et al. (2022)’s DenseNet121 and Kalvankar et al. (2021)’s EfficientNetB5. Some works have investigated combining the advantages of ViT’s attention mechanism with CNNs, such as Wei et al. (2024), Dagli (2023) and Cao et al. (2024). However, application of pure ViTs has been limited and resulted in lower performance than CNNs (Lin et al. 2022), although Kumar et al. (2023) reports fine-tuning to be better than training from scratch. Radio galaxy classification work has been more limited due to lack of large datasets and a well-defined labeling scheme, and so far only employs CNNs; Becker et al. (2021) and Ndung’u et al. (2023) provide comprehensive reviews of the topic. Due to the difficult nature of the data, traditional ML techniques, feature engineering, and finding augmentations specific to radio images, are also active fields of study, while some seek to enhance training datasets with synthetic images using generative deep learning.
The datasets we use to demonstrate the galaxy morphology classification task, GMNIST and RGZ, are both cutouts of single optical or radio galaxies labeled courtesy of the Galaxy Zoo project. Twenty percent of RGZ images contain two labeled galaxies, and four percent contain more than two; these images were all excluded from the classification dataset, reducing the total number of samples to 4692 training images and 1189 test images. Therefore the training data available for RGZ was just over half that available for GMNIST, which was split with 80% of its 10K images in the training set.
Hyperparameters for the classifier training were similar for both datasets, using a learning rate of 0.0005 and a batch size of 64. Only the classifier head was allowed to train, so the backbone is considered frozen in that the weights and biases of the foundation model backbone stay fixed. Classifiers were trained for 100 epochs on GMNIST and 200 on RGZ. Class weights inversely proportional to the number of samples present in the training dataset were used in the cross-entropy loss function used to classify RGZ. To illustrate the performance differences in low- and high-label regimes, the number of samples used for training was varied at 10%, 30%, 50% and 100%, while the test set was kept the same. For comparison, we trained representative models from scratch in a fully-supervised manner in which both the backbone and classifier head were trained. The fully-supervised models were a ViT-Base model with patch size 16, as well as ResNets with both 50 and 18 layers. The best-performing supervised model was ResNet-50, so its results are shown in Figure 3.
3.1 Classification performance
@@
Figure 3 shows classification F1 score as a function of training label percentage. The F1 score is the harmonic mean of precision and recall, calculated using the number of true and false positives for each class, then averaged. While results differ according to dataset, some trends are in common. Performance generally improved when the number of training labels increased, MAE was the worst performer overall, and the smaller ResNet-18 did not classify as well as ResNet-50.
@@
The distillation model AM-RADIO performed almost exactly as well as SigLIP, with a slight advantage in the low-label regime. Both these models demonstrate a F1 score at least  0.2 – 0.25 better than the other models. Self-supervised models DINOv2 and MSN both performed better than the ResNets, although ResNet-50 improved the most when trained on 100% of the data. Unlike DINO and MSN, which use student-teacher paradigms to reconcile partially masked or locally cropped views of images, MAE’s objective is to reconstruct an entire image from a few patches. Balestriero & LeCun (2024) found that the learned representations with the reconstructive power were the least informative for perception. Our findings support that result; it seems that MAE learns features that are irrelevant to classification.
Classification was more difficult with Radio Galaxy Zoo, with F1 scores of less than 0.6 even when using all available training labels. This could be because more of the image is dominated by noise, and less by the galaxy to be classified. The labeling schematic could be another reason for low scores, as discussed in Section 2.1. Only SigLIP and AM-RADIO consistently and significantly out-performed supervised training.
The poor performance on radio galaxy classification is not consistent with results from previous works such as Slijepcevic et al. (2024) and Lastufka et al. (2024), which have shown much higher F1 scores (up to 0.94 for DINOv2) achieved on the MiraBest dataset (Porter & Scaife 2023) using pre-trained foundation models. MiraBest is designed for binary classification in the Fanarhoff-Riley scheme, which could be roughly described as a single Gaussian component with one peak versus a single component with two peaks.
@@
Unlike GMNIST, RGZ images are reconstructions which contain noise with strong inter-pixel correlations on the scale of the synthesized beam. These statistical differences between ImageNet and radio images suggest that a well-performing model may require knowledge of a very specific set of features; a true case of distribution shift.
There are two immediately recognizable visual differences between GMNIST, on which classification worked well, and RGZ (see again Figure 1): radio galaxies occupy a much smaller area of the total image than optical ones, and background noise is much higher and patterned. We experimented with bringing the RGZ dataset into better alignment with GMNIST by leveraging the bounding box information to make very close crops around each radio galaxy. This resulted in a F1 score of 0.70 for all labels, using AM-RADIO to extract frozen features.
While this is not yet on par with classification performance on GMNIST, it is clear that the foundation models are able to extract more information from the closely cropped images. Figure 5 shows the output of selected attention heads on both the cropped (top row) and uncropped (bottom row) RGZ images for two different classes of galaxy. Even though the 1C-1P class of the left side image is in general confidently classified, the attention maps show that areas of high interest to the different attention heads correspond mostly to noise in the uncropped images. The central compact source of the left image only occupies about one $\times×16$ pixel patch in the uncropped image, which has already been upscaled from its original size. Attention maps for sample images from all datasets and all transformer-based backbones can be found in Appendix A. While the out-of-the-box application of foundation models to GMNIST images performed well, it is not sufficient for classification on RGZ where it offers little to no advantage over supervised training, without careful adaptation to suit the chosen foundation model.
Figure 5: Attention maps (red contours) overlaid on closely-cropped (top) and uncropped (bottom) RGZ images. A typical compact source (labeled as 1C-1P) is shown on the left, with a more complex source with diffuse elements (labeled as 2C-3P) on the right. The attention maps are shown for two different attention heads as contours at the 90th percentile level.
3.2 Optimization strategies
We demonstrate various common strategies for improving classification performance on a challenging dataset like RGZ. These include, roughly in order of the time it takes to implement the changes:
  • Use a multi-layer perceptron (MLP) for classification instead of a single linear layer.
  • Use a bigger model with more parameters.
  • Use different pre-processing methods.
  • Fine-tune the backbone via Low-Rank Adaptation (LoRA).
  • Fine-tune the backbone using a semi-supervised method.
The first two strategies consist of adapting the model, the third adapts the data to suit the model, and the last two are different ways to implement fine-tuning. For purposes of illustration, we report the results using all available training labels and the models AM-RADIO and ResNet-50 in Table 4. With the notable exceptions of employing a close crop and fine-tuning, these techniques improved the classification result of AM-RADIO more than that of ResNet-50. In fact, using either larger input image sizes or larger networks decreased the accuracy of ResNet.
The effects of each technique is reported on its own; one can expect that employing multiple techniques is a little less than the sum of each individual improvement. For example, using AM-RADIO and combining a close crop and fine-tuning the backbone gives an F1 of 0.82, rather than 0.84. We briefly describe each technique and its impact on classification score below.
Technique ResNet-50 F1 AM-RADIO F1
Whitening Layer 0.519 (+0.008) 0.603 (+0.013)
MLP 0.526 (+0.015) 0.623 (+0.034)
RN101/ViT-Large 0.486 (-0.025) 0.625 (+0.036)
RN152/ViT-Huge 0.486 (-0.026) 0.653 (+0.063)
Crop 0.643 (+0.132) 0.705 (+0.115)
Upsample 320 0.481 (-0.031) 0.630 (+0.041)
Upsample 480 0.465 (-0.046) 0.666 (+0.080)
Upsample 512 0.454 (-0.057) 0.667 (+0.078)
LoRA 0.674 (+0.163) 0.692 (+0.102)
Fine-tuning 0.737 (+0.226) 0.721 (+0.132)
Table 4: Effect of various optimization techniques on RGZ classification.
3.2.1 Multi-layer perceptron
So far, classification was done using a single linear layer as a classifier. Increasing the number of fully-connected layers in the classifier head lends a few advantages. Mainly, the addition of at least one hidden layer enables the classifier to learn non-linear functions. These dense networks with one or more hidden layers are commonly referred to as MLPs (multi-layered perceptrons).
For a two-layer MLP, we observe a 5.7% increase in the F1 score of AM-RADIO and a 2.9% increase for ResNet-50.
We also investigate the use of a whitening layer (Kalapos & Gyires-Tóth 2024), which scales the frozen features before input to the linear classifier layer. Effectively, this is a two-layer MLP with the weights and biases of the input layer pre-calculated. This leads to a 2.2% increase in the F1 score of AM-RADIO and a 1.5% increase for ResNet-50.
3.2.2 Bigger model
In general, performance increases with the number of parameters in a model (e.g. Zhai et al. 2022). AM-RADIO has pre-trained two models larger than 86M parameter ViT-Base: ViT-Large with 307M parameters, and ViT-Huge with 632M parameters. Larger ResNet variants are ResNet-101, with 42.8M parameters, and ResNet-152, with 58.5M parameters, all pre-trained for ImageNet classification.
Using features extracted from both these models, we observe an increase in the F1 score of AM-RADIO of 8.1% (ViT-Large) and 10.7% (ViT-Large). ResNets with more parameters did not lead to an improvement, but rather the opposite; the F1 score decreased by 5% when using features extracted from ResNet-101, and by about the same amount for ResNet-152.
3.2.3 Custom data pre-processing
Like most scientific images, optical and radio astronomy data is not delivered straight from the telescopes. Raw data must be calibrated, to account for instrumental or environmental effects as much as possible, and sometimes further processed like in the case of imaging the visibility data collected by radio arrays. Much more so than the natural image domain, the data reduction pipeline has a large impact on the capabilities of the resulting dataset.
In the case of RGZ images, a number of steps have taken place before the final images are incorporated into the dataset. First, radio sources were identified from a survey catalog, which is itself the result of a source-finding algorithm. Next, various filters such as signal-to-noise ratio determined if the image source was kept in the sample or not. Were archival images not available, the initial image would have had to be reconstructed. Experts then had to decide on a cutout size and how to scale each image, which consist of float values spanning five or more orders of magnitude, before finally converting the image to a PNG. Often, both image reconstruction and scaling favor a certain type of radio emission over another, making a dataset more suited to certain scientific goals or downstream tasks than others.
The final product of most astrophysics image reduction pipelines are 16- to 32-bit FITS images in physical units, rather than 8-bit PNG scaled between 0 and 255. As only PNG images were available in the GMNIST and RGZ datasets, we can only examine pre-processing methods that operate on those images, which have already lost some detail. One way in which the data has been adapted to fit the model was already mentioned previously; the images had to be re-sized in order to be accepted as input to most models with a ViT-Base backbone. CNNs do not have this input size requirement of $\times×224$ pixels, as the convolution kernel can operate on images of any size and shape. Similarly, AM-RADIO was designed to take any input image size as long as the dimensions are a multiple of 14.
We resize the RGZ images to 320, 480, and 512 pixels square using bilinear interpolation. Interestingly, increasing the image size does not necessarily increase performance; the optimal size for AM-RADIO is $\times×480$, leading to a F1 score increase of 8%. As with increasing the model size, increasing image size has a detrimental effect on the performance of classification via ResNet-50.
Another pre-processing method already mentioned is making a close cutout of the radio source based on the bounding box information. Cropping to the area around the central galaxy leads to a large improvement in F1 score: 16% increase for AM-RADIO, and 26% increase for ResNet-50. Similar pre-processing, involving cropping images to less than 50% their original area, was used since Dieleman et al. (2015) and Zhu et al. (2019) to classify optical galaxy images from SDSS and DeCALs dataset; more recent works like Wei et al. (2024) and Urechiatu & Frincu (2024) also use this method to achieve F1 scores above 0.95. As seen from the attention maps in Figure 5, the close crop means that more patches of the ViT contain parts of the object to be classified, allowing the network to use information from more than just one or two patches to perform classification. Custom pre-processing allows the data to complement features the foundation model has already learned, and emphasizes the importance of domain-specific expert knowledge in the development of machine learning pipelines.
3.2.4 Full fine-tuning of the backbone model
This is traditional fine-tuning: unfreezing parameters the backbone network and allowing them to update. It can be done layer-by-layer or block-by-block, but for this experiment we simply unfreeze the entire backbone after ten epochs, which are used to allow the newly initialized classifier layer to acclimate.
When fine-tuning the backbone, we observe an increase in the F1 score of AM-RADIO of 22% and 44% for ResNet-50. This is the largest improvement in performance by any of the optimization techniques we investigated.
3.2.5 Fine-tuning the backbone model using LoRA
Low-Rank Adaptation, or LoRA, is a parameter-efficient fine-tuning method. Instead of directly updating model weights, LoRA tracks changes to the weights and then decomposes the large matrix of weight changes into smaller matrices. These low-rank matrices are multiplied to approximate the full matrix of weight changes, allowing for fast and efficient fine-tuning. In practice, it is not much more difficult to implement than full fine-tuning, although there are additional hyperparameters to be optimized.
Using LoRA leads to a 17.3% increase in the F1-score with AM-RADIO as a backbone, and a 31.8% increase when using ResNet-50. These are substantial improvements that result from tuning less than 5% of the model parameters.
@@ 4 Source detection
Source detection in both optical and radio images is largely performed analytically. Work using deep learning to replace these time-consuming analytic pipelines has been done by Wu et al. (2019); Farias et al. (2020); Burke et al. (2019); Riggi et al. (2023), mainly through application of Faster- or Mask-RCNN. Sortino et al. (2023) provides an overview of both CNN and ViT-based methods, including YOLO. More recent results based on simulated radio data include Taran et al. (2023) and Cornu et al. (2024), which both surpass the capabilities of analytic source detection.
Here we performed source detection on radio continuum datasets RGZ, which mostly contains single galaxies labeled by morphology, and MGCLS, where a single image contains from 10-100 individual compact sources. MGCLS images may also contain larger sources which are not labeled.
@@
Table 5: Source detection mAP@50 for the RGZ and MGCLS datasets. Results for frozen and fine-tuned backbones are compared with supervised baselines.
@@
As was done for classification, images were upscaled from their native resolution. For RGZ, scaling to $\times×224$ meant that the average single object size became $\times×55$ pixels, while the 25th percentile was $\times×24$ pixels. MGCLS crops were doubled in size from $\times×256$. This was done so that the average compact source occupied $\times×16$ pixels, and the 25th percentile $\times×14$ pixels. The size of the sources relative to the transformer patch size is important; sources unable to occupy most of a patch were difficult to detect. This was proven empirically – performance on detection with MGCLS increased by more than 10% when the image size doubled. Otherwise, we kept the default training data augmentations: a randomly applied blur, contrast adjustment, and color jitter.
@@
Our first experiment kept the backbones frozen, allowing only Faster-RCNN’s detection head layers (FPN, RPN and ROI) to train. Due to differences in architecture this resulted in a different number of trainable parameters: 21M for ViT-Det, which was used for all the ViT-based models, 17.8M for ResNet-50 with FPN, and 11M for ResNet-18. Because of the structure of the FPN for ViT-Det, for this task the DINO backbone used was DINOv1, which had available pre-trained weights for the architecture ViT-Base with 16x16 patch size (DINOv2 uses 14x14). In the second experiment we allowed the weights and biases of the backbone networks to update as well, in the process known as fine-tuning.
Generally, a learning rate of 0.001 was used, with the exception of training ViT-Base on MGCLS from scratch, and both frozen and fine-tuned MAE on MGCLS, where we used a learning rate of 0.0005 in order to ensure a smooth decay of the training loss. Training was for 100 epochs, with a batch size of 16, performed on 1-2 GPUs depending on availability. Due to training requiring a number of hours to complete, we did not repeat training runs in order to report a range in our chosen metric. we report the mean average precision at an intersection-over-union (IOU) threshold of 50%, known as the mAP@50, in Table 5. This metric considers a prediction correct if the overlap between the predicted and actual object bounding boxes is at least 50%, and is commonly used to evaluate object detection tasks.
Results showed that MSN and DINO are the best foundation models for this task. They outperformed training from scratch, even when the backbone was frozen, except for MSN on MGCLS. Transformers pre-trained using multi-modal datasets (SigLIP and AM-RADIO) performed worse than the other ViTs, including MAE, on RGZ, and slightly better on MGCLS, although not as good as DINO and MSN. This is an interesting result, as both SigLIP and AM-RADIO were the best at classification. Representations learned through text-image pairs appear to be less relevant to object detection in images than they are to classification.
@@
Although the ResNet models did not outperform training from scratch when used as a frozen backbone, they equaled or exceeded that performance when fine-tuned. Detection on galaxies in the RGZ dataset appears to be a task particularly suited to ResNets, while Vision Transformers do very well at detecting compact sources in MGCLS. Generally, detection was very good on MGCLS for networks larger than ResNet-18. Possible reasons for the higher mAP@50 include the single category (compact source) and the tens of sources present in the images compared to the 1-6 sources in RGZ. Additionally, bounding boxes in RGZ can be quite large, including a lot of noise and making them harder to predict accurately.
@@
These results show that transfer learning, starting with the pre-trained weights of a foundation model, is a good practice, as the right choice of model leads to better results even using a frozen backbone than training from scratch, and fine-tuning brings further improvements. It is important to ensure that properties of the architecture, such as patch size in the case of ViTs, are complementary to the data being used for the downstream task.
5 Best Practices Guidelines
We do not claim that the the two datasets and two tasks examined here are completely representative of the many applications of deep learning for astrophysics. The lessons learned from considering various methods in using foundation models in these circumstances may be useful to practitioners, so we gather the results here into a general guideline.
As always, the dataset type – characteristics of the images and type and number of available labels – has the most impact on DST performance. These properties will inform the choice of foundation model, which in turn will determine how the dataset should be customized to complement the model itself. While GMNIST images are complementary enough to ImageNet and similar pre-trained backbones, a dataset like RGZ might benefit from a model with different pre-training. Models pre-trained on images with strong background patterns and textures, such as satellite or thermal imagery, or perhaps images from medical scans or microscopy could be good avenues to investigate. The labeling scheme of the dataset is also something to consider; the RGZ labels, while sensible from a scientific point of view, is not well suited to machine learning, as information about the number of radio components is obscured during scaling and conversion of the original images to PNG format.
Besides selection of the foundation model for the backbone, there is also the choice of projector head for the downstream task. Clearly, some projector heads will have to be specifically designed for the task, such as Faster-RCNN for object detection. Other possibilities for detection and segmentation include YOLO, which is one of the best performing architectures on natural images, and can be used with a ResNet backbone. From a practical perspective, the choice of projector might also limit the choice of backbone; YOLO’s latest versions can be adapted for ResNet backbones but not yet for ViTs. For simpler tasks like classification, the projector head is less important; we found that modifying it with an extra hidden layer or whitening layer only brought improvements of at most 5%.
Fine-tuning the model, once selected, yields the most significant improvements. Model and dataset size determine the necessary time and resources; with many astrophysics datasets having less than 10K images, fine-tuning a model of a few million parameters should be feasible. An alternative to full fine-tuning is to only un-freeze the upper layers or blocks of a network. If computational resources are very limited, LoRA is an alternative that will generally be a few percent less accurate than full fine-tuning.
If fine-tuning does not reach the required performance metric, or if one wishes to generally improve the system as much as possible, one should carefully re-consider the data pre-processing. Together with custom dataset assembly and labeling, this is the most labor intensive step and requires the most specialized domain knowledge – however it will also have the biggest impact on performance.
For example, the simple task of resizing images to match the input dimensions of the model is often necessary, but can be done optimally. We saw that resizing images can boost performance by up to 8%, but that bigger was not necessarily better and was in fact worse when using ResNet-50. Furthermore, our example of RGZ classification showed that cropping images to focus on regions of interest was the next-best thing to fine-tuning. Considering the earlier discussion of the RGZ labeling scheme in Section 3.1, we also speculate that using different scalings of the original FITS images to include information about not only bright peaks but also flux islands could significantly improve distinction between visually similar classes like 2C-2P and 2C-3P (see Figure 1 for example). One technique seen in the literature (Aniyan & Thorat 2017, used by Maslej-Krešňáková et al. (2021) for example) is sigma-clipping, where noise is removed by setting all pixels below a certain threshold to zero.
Note that in this work we do not discuss data augmentation, which is normally done during training to help the network learn more robust representations. Several works address augmentation schemes for both optical and radio images; Hayat et al. (2021) appendix D for optical, and Slijepcevic et al. (2024) section 3.9 or Maslej-Krešňáková et al. (2021) section 4.2 for radio.
These general principles can help guide effective application of vision foundation models to astrophysical image data. Experimentation is unlikely to be a linear process; practitioners may need to iteratively revisit choices of foundation models, dataset preparation, and task-specific projector heads in order to achieve the desired result.